# Integrating Random Regret Minimization-Based Discrete Choice Models with Mixed Integer Linear Programming for Revenue Optimization

### v2

@@ 3.1 Problem definition
Random regret minimization asserts that individuals choose alternatives to minimize the regret they may experience from foregone options. Randomness arises from the inability of researchers to observe all the regret a customer perceives. To illustrate the RRM mechanism, consider the example of passenger travel mode choice provided by chorus2008random . The alternatives are car (1), train (2), and bus (3), each with attributes $x_{1},\ x_{2}$ and $x_{3}$ representing time, cost, and alternative-specific dummy variables of travel. The regret for each alternative is calculated, and a rational customer would choose the alternative with the minimum regret:
$R_{i}=\sum_{j\neq i}R_{ij}\ \text{and}\ R_{ij}=\sum_{k\in\{x_{1},x_{2},x_{3}\}}R_{ijk},\ \text{and}\ R_{ijk}=\max\{v_{ok},\beta_{k}(k_{j}-k_{i})+v_{k}\}$
Where $\beta$ parameters, representing consumers’ tastes for each attribute $k$, are estimated from historical data, and $k_{i}$ represents the $k^{th}$ attribute of alternative $i$. Regret, partially known to researchers, is modified by adding an error term, drawn from an independent and identically distributed (iid) extreme value type I distribution, to obtain the overall regret and $v_{ok}$ and $v_{k}$ stand for unobserved regret in attribute perceptions chorus2010new . Additionally, let $I=|C|$ represent the number of alternatives in the universal choice set $C$, indexed by $i$, where $i=0$ denotes the opt-out option when the customer chooses not to buy any alternatives. The set $N$ represents customers, indexed by $n$. For each customer, the choice set $C_{n}$ is considered, with each product $i$ available in quantity $c_{i}$. The supplier aims to maximize revenue by making decisions on whether a product $i$ should be offered to customer $n$ and what price of alternative $i$ to be offered to customer $n$, denoted as $p_{in}$. Customers are served sequentially based on a priority list, determined by factors such as customer loyalty, arrival order, customer segments, anticipated order quantities, etc. Such an assumption is standard (e.g., see pacheco2016new ; PACHECOPANEQUE202126 ). As it was shown by pacheco2016new ; paneque2017integrating , when solving this problem directly using DCMs, highly non-linear and non-convex problems arise. Therefore, to address this challenge, we employ MILP models. This paper explores two cases, the first assumes the supplier has no capacity limitation on supplies, while the second considers a scenario where there is a capacity limitation on the supply.
@@ 3.2.1 Sets
$I=|C|$ Number of alternatives, indexed by $i$ ($i=0$ denotes the no-purchase
alternative or opt-out option)
$L_{in}$ Number of binary variables characterizing the price levels for each alternative
and individual, indexed by $I$
$N$ Number of customers, indexed by $n$
$R$ Number of draws, indexed by $r$
@@ 3.2.2 Input Parameters
$c_{i}$ Available quantity of alternative $\text{Available quantity of alternative }i\text{ for sale, maximum capacity of service }i$
@@
$pm$ Price multiplier, to get desired price levels. Also, to set the price levels and
ranges precisely, one can increase or decrease the value of the multiplier
@@
with alternative $j$ for customer $n$ and scenario $r$. Note that:
$ER_{ijnr}=\sum_{m=1}max\{v_{onmr},\beta_{nm}(x_{jm}-x_{im})+v_{xnmr}\}$
where $v_{onmr},v_{xnmr}$ are draws of Gumbel$(0,1)$. $x_{jm}$ represents the $m^{th}$ level of
alternative $j$’s attribute $x$. It is worth mentioning that price related regret has been
excluded from $ER_{ijnr}$
$mm_{ijnr}$ $=\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}\ \text{are upper bounds on variables $RR_{ijnr}$}$
$ll_{ijnr}$ $=\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}\ \text{are lower bound on variables $RR_{ijnr}$}$
@@
$m_{inr}$ $=\sum_{j\neq i}\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}+ER_{ijnr},\ \text{upper bounds on variables}\ R_{inr}$
$l_{inr}$ $=\sum_{j\neq i}\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}+ER_{ijnr},\ \text{lower bounds on variables}\ R_{inr}$
$m_{nr}$ $=\max_{i\in C_{n}}\{m_{inr}\}\ \text{upper bounds on variables }R_{nr}$
$l_{nr}$ $=\min_{i\in C_{n}}\{l_{inr}\}\ \text{lower bounds on variables }R_{nr}$
[…]
$\beta_{nm}$ Customer $\text{Customer }n\text{ taste for attribute }m\text{ of alternatives}$
@@ 3.2.3 Variables
$b_{ijnr}\in\{0,1\}$ Binary variables to capture the maximum of $RR_{ijnr}$
@@ 3.2.4 Uncapacitated supply MILP model
$\max\frac{1}{R}\sum_{n}\sum_{i>0,i\in C_{n}}\sum_{r}(lp_{in}w_{inr}+\sum_{l}l\alpha_{inrl}pm)$ (1)
@@
There are $R$ scenarios for each customer in the objective function 1, and the average revenue will be computed by dividing the total revenue over $R$, and $lp_{in}$ is a lower bound on the price levels. For an alternative $i$ and customer $n$, exactly one price level should be offered in scenario $r$. If alternative $i$ is chosen ($w_{inr}=1$), the price level should be selected ($\lambda_{inl}=1$) in a way that $lp_{in}+\sum_{l}l\alpha_{inrl}pm$ has the most positive contribution to the objective function while respecting constraints of the model. We define $pm$ as a constant price multiplier, and integers $l$ to characterize price levels. The opt-out alternative is excluded from the objective function as it does not affect the revenue. Constraints 2,3, 4, and 5 are to calculate the anticipated regret arising from the comparison of the price attribute of two alternatives. In this model, all alternatives are available for each customer. An error term $v_{nr}$ has been added to the price-raised regret between alternatives $i$ and $j$, $RR_{ijnr}$, to unveil the maximum anticipated regret.
The parameters $B_{np}$ shows customer $n$ taste for the price attribute, assumed to be negative for all customers. However, considering the RUM-based MILP, we do not compare alternatives pairwise to calculate regret. Instead, we measure the regret obtained from different price levels for each alternative. Then, the alternative rendering the minimum regret will be chosen. This is the major difference between the RUM-based and RRM-based MILPs.
The terms $ER_{ijnr}$ are the sum of the anticipated regret of attributes (other than the price attribute) for alternatives, and constraints 6 are to obtain the overall anticipated regret of an alternative using these terms and price attribute regret.
Constraints 7 and 8 determine the alternative with the lowest regret among all other alternatives in the choice set of a customer at each scenario. Here, the alternative providing the least regret will be chosen, while in RUM-based models, the one with the highest utility is chosen.
Constraints 9, 10 and 11 have been used to linearize the objective function. Constraints 12 illustrates the price bounds for each single customer at each scenario. $mp_{in}$ stands for upper bounds on the price, and $lp_{0n}=mp_{0n}=0$.
@@
The rest of the constraints are bound constraints where the bounds were defined in the notation part:
@@
These constraints typically enhance the efficiency of the model-solving process.

### v3

@@ 3.1 Problem definition
RRM posits that individuals select alternatives to minimize the regret they may encounter from foregoing other options. In other words, they conduct pairwise comparisons of the regret associated with each alternative or product, including the option of not choosing any alternative (opt-out), and ultimately choose the alternative expected to result in the least regret. Randomness in the context of RRM arises from researchers’ inability to observe all the regret a customer experiences. To elucidate the RRM mechanism, consider the example of passenger travel mode choice provided by [9]. In this scenario, three alternatives are considered: car, train, and bus, each characterized by attributes $k^{1}$, $k^{2}$, and $k^{3}$, representing time, cost, and alternative-specific dummy variables of travel such as comfort level, reliability, accessibility, or any other characteristic that distinguishes one mode of transportation from another. The regret for each alternative is calculated as follows, and a rational customer would opt for the alternative with the minimum regret:
$R_{i}=\sum_{j\neq i}R_{ij}\ \text{and}\ R_{ij}=\sum_{k\in\{k^{1},k^{2},k^{3}\}}R_{ijk},\ \text{and}\ R_{ijk}=\max\{v_{ok},\beta_{k}(k_{j}-k_{i})+v_{k}\}$
Where $\beta$ parameters, representing consumers’ tastes for each attribute $k$, are estimated from historical data, and $k_{i}$ represents the $k^{th}$ attribute of alternative $i$. Regret, partially known to researchers, is modified by adding an error term, drawn from an independent and identically distributed (iid) extreme value type I distribution such as Gumbel distribution, to obtain the overall regret and $v_{ok}$ and $v_{k}$ stand for unobserved regret in attribute perceptions [10]. Thus, given two alternatives $i$ and $j$, we first compute the regret of choosing alternative $i$ attribute-wise, i.e., $R_{ijk}$, and then we obtain the sum of attribute-level regret, i.e., $R_{ij}$. Finally, by summing all the regret caused by selecting alternative $i$ given all other alternatives, we obtain the total anticipated regret caused by selecting alternative $i$, i.e., $R_{i}$. We will compute these values in MILP for each customer and all available alternatives for each customer. Additionally, let $I=|C|$ represent the number of alternatives in the universal choice set $C$, The set $N=\{1,...,n\}$ represents customers. For each customer $n$, the choice set $C_{n}$ is considered, with each product $i$ available in quantity $c_{i}$. Note that $i=0$ shows the opt-out alternative. The supplier aims to maximize revenue by making decisions on whether a product $i$ should be offered to customer $n$ and what price of alternative $i$ to be offered to customer $n$, denoted as $p_{in}$. Customers are served sequentially based on a priority list, determined by factors such as customer loyalty, arrival order, customer segments, anticipated order quantities, etc. Such an assumption is standard (e.g., see [2, 52]). As it was shown by [2, 16], when solving this problem directly using DCMs, highly non-linear and non-convex problems arise. Therefore, to address this challenge, we employ MILP models. This paper explores two cases, the first assumes the supplier has no capacity limitation on supplies, while the second considers a scenario where there is a capacity limitation on the supply.
@@ 3.2.1 Sets
$C$ Set of all alternatives. $C_{n}$ is the set of all available alternatives for customer $n$
$L_{in}$ Set of price levels for each alternative $i$ and customer $n$ indexed by $I$
$N$ Set of customers, i.e., N= {1,…,n}
$Ra$ Set of random draws, where each draw is indexed by $r$ and $Ra=\{1,..,r\}$
Note that $C_{n}$ is of significance when the model is capacitated, as prior customers to customer $n$, for instance, can purchase all available quantity of an alternative. We assume that there is an order list from which the supplier satisfies the orders of the customers accordingly.To give an example of $L_{in}$, suppose $L_{in}=\{1,2,3\}$. Additionally, there is a fixed parameter for pricing, $pm$, let us assume $pm=0.5$. Then, the price levels of alternative $i$ for customer $n$ will be $0.5,1,$ and $1.5$, which are obtained by element-wise multiplication of $pm$ by the elements of $L_{in}$. $pm$ allows us to adjust the price of alternatives for customers. Finally, $Ra$, represents the set of scenarios or draws. This accounts for the stochastic behavior of customers, and we solve our models for several draws to obtain robust results.
@@ 3.2.2 Input Parameters
$c_{i}$ Available quantity of alternative $\text{Available quantity of alternative }i\text{ for sale, maximum capacity of the alternative }i$
@@
alternative is zero.
$pm$ The price multiplier is used to attain desired price levels. By adjusting the value
of the multiplier, one can precisely set the price levels and ranges. Increasing or
decreasing the multiplier value allows for fine-tuning of prices.
@@
with alternative $j$ for customer $n$ at scenario $r$. Note that:
$ER_{ijnr}=\sum_{x}\sum_{k\in\{k_{1},...,k_{a}\}}max\{v_{onxr},\beta_{nx}(x_{kj}-x_{ki})+v_{nxkr}\}$
where $v_{onxr},v_{nxkr}$ are draws of Gumbel$(0,1)$ for customer $n$, and $k^{th}$ level of
attribute $x$ at draw $k$. Note that $x_{kj}$ represents the $k^{th}$ level of alternative $j$’s
attribute $x$. Note that the price attribute has been excluded from $ER_{ijnr}$
since it is a variable in our models.
$mm_{ijnr}$ Upper bounds on variables $RR_{ijnr}$ equivalent to $\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}\}$
$ll_{ijnr}$ Lower bounds on variables $RR_{ijnr}$ equivalent to $\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}$
@@
$m_{inr}$ Upper bounds on variables $R_{inr}$, which is $\sum_{j\neq i}\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}+ER_{ijnr}$
$l_{inr}$ Lower bounds on variables $R_{inr}$, which is $\sum_{j\neq i}\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}+ER_{ijnr}$
$m_{nr}$ Upper bounds on variables $\text{Upper bounds on variables }R_{nr},\textbf{equivalent to}\max_{i\in C_{n}}\{m_{inr}\}$
$l_{nr}$ Lower bounds on variables $R_{nr}$ which is equivalent to $\min_{i\in C_{n}}\{l_{inr}\}$
$M_{nr}$ Big-M notion, equivalent to $m_{nr}-l_{nr}$
$M$ Big-M notion, equivalent to $\max_{\forall i,j\in C_{n},i\neq j,n,r}\{|ll_{ijnr}+ER_{ijnr}|,|mm_{ijnr}+ER_{ijnr}|\}$
$\beta_{np}$ Customer $n^{th}$ taste for price attribute of alternatives.
$\beta_{nm}$ Customer $n^{th}$ taste for attribute $m$ of alternatives.
Note that all upper bound and lower bound parameters aid in potentially solving the MILP efficiently and quickly. However, their inclusion in the model is not mandatory. Specifically, setting upper bounds on Big-M, which assists in activating and deactivating the constraints of the MILP models, usually enhances solution time. Regarding $\beta_{np}$, it indicates customer $n$’s preference for the price attribute. For instance, if $\beta_{np}=-1$, and the customer purchases the alternative, it implies that the customer is willing to pay one dollar for each unit of the alternative.
@@ 3.2.3 Variables
$b_{ijnr}\in\{0,1\}$ Binary variables to capture the $R_{inr}$
@@
Among the variables, $y_{in}$ relates to the direct decision of the supplier to grant access to alternative $i$ for customer $n$ or not. However, $y_{inr}$ relates to the availability of alternative $i$ for customer $n$ in scenario $r$. In other words, it not only considers the supplier’s decision regarding the availability of alternative $i$ for customer $n$, but also takes into account the available capacity of alternative $i$ through the constraints of the MILP model. Variables $\alpha_{inrl}$ are technical variables used to linearize the objective function of the MILP models presented in this paper. When $\alpha_{inrl}=1$, it indicates that customer $n$ purchases alternative $i$ in scenario $r$ at price level $l$; otherwise, it is equal to zero. Similarly, variables $b_{ijnr}$ are related to techniques in MILP models to activate or deactivate constraints. In our paper, these variables assist in computing the anticipated regret of purchasing alternative $i$ by customer $n$ in scenario $r$. Variables $z_{ijnr}$ assist us in computing the discounted regret of choosing alternative $i$ over alternative $j$ by customer $n$ in scenario $r$. We refer to it as ”discounted” following Pacheco et al. (2016) [2], as the purchasing decision has not yet been finalized by the customer, and we are anticipating the regret. The remaining variables pertain to the calculation of regret and the selection of the final choice.
@@ 3.2.4 Uncapacitated supply MILP model
The relationships between variables are captured through the MILP models. In this part, we first present our models, and then we will explain the constraints and relationships between variables.
$\max\frac{1}{|Ra|}\sum_{n}\sum_{i>0,i\in C_{n}}\sum_{r}(lp_{in}w_{inr}+\sum_{l}l\alpha_{inrl}pm)$ (1)
@@
There are $|Ra|$ scenarios for each customer in the objective function 1. This is due to the stochastic nature of consumer behavior. Averaging over several scenarios will result in more robust solutions. The $lp_{in}$ variable represents a lower bound on the price levels, i.e., the minimum price at which alternative $i$ will be offered to customer $n$. Additionally, exactly one price level should be selected in each scenario $r$, which maximizes revenue while respecting the constraints of the model, i.e., minimizing customer regret. If alternative $i$ is chosen (i.e., $w_{inr}=1$) for the customer $n$ in scenario $r$, the corresponding price level ($\lambda_{inl}=1$) should be selected in such a way that $lp_{in}+\sum_{l}l\alpha_{inrl}pm$ contributes most positively to the objective function, while still respecting the constraints of the model, as it is a maximization problem. We define $pm$ as a constant price multiplier, and integers $l$ are used to characterize price levels. The opt-out alternative is excluded from the objective function, (i.e., $i>0$) as it does not affect revenue (i.e., the price of the opt-out alternative is zero).
To calculate the anticipated regret arising from the comparison of the price attribute of two alternatives, we introduce constraints 2,3, 4, and 5. In the uncapacitated version, the above MILP model, all alternatives are available for each customer. An error term $v_{nr}$ has been added to the price-raised regret between alternatives $i$ and $j$, $RR_{ijnr}$. Finally, we compare this anticipated regret with that of unobserved heterogeneity, $v_{onr}$, to determine the maximum anticipated regret. In more details,
$RR_{ijnr}=\max\{v_{onr},\beta_{np}(lp_{jn}+\sum_{l}l\lambda_{jnl}pm-lp_{in}-\sum_{l}l\lambda_{inl}pm)+v_{nr}\},\quad\forall i,j\in C_{n},i\neq j,n,r,$
where $lp_{jn}+\sum_{l}l\lambda_{jnl}pm$ characterizes the price of the alternative $j$ for customer $n$. Since $RR_{ijnr}$ is the maximum of two terms, it can be explicitly asserted that:
$v_{onr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$
$\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr}\leq RR_{ijnr},\quad$ $i,j\in C_{n},i\neq j,n,r$
However, $RR_{ijnr}$ is exactly equal to the maximum of these two terms, which is ensured by constraints 3 and 5 as follows:
If $b_{ijnr}=1,\forall i,j\in C_{n},i\neq j,n,r$, it is obvious that:
$RR_{ijnr}\leq v_{onr}+M_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r$
These constraints will be inactive since:
$ll_{ijnr}\leq v_{onr},$ $\forall i,j\in C_{n},i\neq j,n,r$
$mm_{ijnr}-ll_{ijnr}=M_{ijnr}$
$RR_{ijnr}\leq mm_{ijnr}\leq v_{onr}+mm_{ijnr}-ll_{ijnr}=v_{onr}+M_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r$
meaning that constraints 3 always hold if $b_{ijnr}=1,\forall i,j\in C_{n},i\neq j,n,r$. At the same time, due to constraints 5, when $b_{ijnr}=1,\forall i,j\in C_{n},i\neq j,n,r$,
$RR_{ijnr}\leq\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr},\quad\forall i,j\in C_{n},i\neq j,n,r$
Due to constraints 4, we also have:
$\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r$
This implies that
$\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr}=RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r$
when
$v_{onr}\leq\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr}$
With the same rationale, when $b_{ijnr}=0$, we have $v_{onr}=RR_{ijnr}$. Indeed, these sets of constraints help us to linearize
$RR_{ijnr}=\max\{v_{onr},\beta_{np}(lp_{jn}+\sum_{l}lp_{jn}\lambda_{jnl}-lp_{in}-\sum_{l}lp_{in}\lambda_{inl})+v_{nr}\}$
Considering the RUM-based MILP, we compare alternatives pairwise to calculate regret. We compute the regret obtained from different price levels for each alternative pairwise, i.e., in comparison to other alternatives. Then, the alternative rendering the overall minimum regret will be chosen. The parameters $ER_{ijnr}$ represent the sum of the anticipated regret of attributes (other than the price attribute) for alternatives, and constraints 6 are used to obtain the overall anticipated regret of an alternative by summing $ER_{ijnr}$ and the regret related to the price attribute, i.e., $RR_{ijnr}$. Note that $ER_{ijnr}$ are parameters of the model and they have been calculated beforehand.
It is very unlikely to have two alternatives with the minimum regret. Without loss of generality, we assume that ties are not allowed. Constraints 6, 7 and 8 in our model, capture the alternative with the lowest regret as follows:
$R_{nr}\leq R_{inr},$ $\forall i\in C_{n},n,r,$ (14)
$R_{inr}-M_{nr}(1-w_{inr})\leq R_{nr},$ $\forall i\in C_{n},n,r$ (15)
Variables $R_{nr}$ capture the minimum regret among all alternatives in the choice set of customer $n$ at scenario $r$. The overall regret of alternative $i$ for the customer $n$ at scenario $r$ is measured by variable $R_{inr}$ through constraint (13). Let us assume that $i$ is the alternative with the minimum regret at scenario $r$ for customer $n$ in her choice set. So, $w_{inr}=1$, i.e., the customer $n$’s choice at scenario $r$ is alternative $i$ as it brings about the least anticipated regret among alternatives. Due to constraints (14) and (15), by substituting $w_{inr}=1$, we have
$R_{nr}\leq R_{inr}\quad\text{and}\quad R_{inr}\leq R_{nr}$
Which can be inferred that,
$R_{nr}=R_{inr}$
There is not any other alternative such as $j$ that $w_{jnr}=1$. Otherwise, $R_{inr}\leq R_{nr}$ and $R_{jnr}\leq R_{nr}$. Moreover, we have $R_{nr}\leq R_{inr}$ and $R_{nr}\leq R_{jnr}$. We can conclude that $R_{nr}=R_{inr}=R_{jnr}$. This is in contradiction with the assumption that ties are not allowed. Hence, there exists exactly one alternative such as $i$ for which $w_{inr}=1$ and $R_{nr}=R_{inr}$. Hence, $w_{jnr}=0,\forall j\neq i,j\in C_{n}$. It can be checked that constraints (13), (14), and (15) are inactive when $w_{jnr}=0$. So far, we have shown how the alternative with the minimum anticipated regret will be selected by a customer in different scenarios. However, note that we have assumed that the prices of the alternatives have been chosen by the other constraints. In other words, there are intricate relationships between pricing and the choice of the customers. We have just covered how the anticipated regret of alternatives has been calculated. Part of this regret comes from the price of the product, which affects constraints (2), (3), (4), and (5) by influencing the $RR_{ijnr}$ through variables $\lambda_{inl}$ in constraints 3. For each alternative $i$ and customer $n$, this binary variable will take a value of 1 for only one price level, i.e., $\sum_{l}\lambda_{inl}=1$, as indicated by constraints (11). Due to these constraints and since it is a maximization problem, it is certain that $\lambda_{inl}$ and $w_{inr}$ will each be equal to exactly one for an alternative $i$ at scenario $r$ and a price level $l$, respecting other constraints of the model. Thus, constraints (9) and (10) will take the form $\alpha_{inrl}\leq 1$. Again, since it is a maximization problem, $\alpha_{inrl}=1$ so that it has the most possible positive contribution to the objective function. This is how constraints help us compute the regrets and select the alternative with the lowest anticipated regret while maximizing the revenue and pricing the products. Note that the constraints (11) help us establish an upper bound for the price of a specific alternative $i$ and customer $n$ by not allowing the chosen price to exceed the threshold, i.e., $mp_{in}$. This is useful when categorizing our customers; for example, the product price may be cheaper for students.
@@
The remaining constraints are bound constraints, where the bounds were defined in the notation part.
@@
These bounds do not affect the quality of the solutions; however, these constraints typically enhance the efficiency of the model-solving process.